OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition.

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. passes through (1, –20) and (18, 3)

OpenStudy (anonymous):

First try to find the slope of the line. Can u?

OpenStudy (anonymous):

(y2 -y1) SLOPE = --------- (x2-x1)

OpenStudy (anonymous):

i have 23/17

OpenStudy (anonymous):

good

OpenStudy (anonymous):

Now, the equation of the line is (y-y1) =m (x-x1)

OpenStudy (anonymous):

where m is the slope of the line

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

yes thanks

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

AND remember to take in the slope-intercept form which is y=mx+c